OpenStudy (anonymous):

the sum of three integral numbers is 117 find the numbers.

OpenStudy (anonymous):

3x + (x+1) + (x +2) = 117?.... right

OpenStudy (anonymous):

another good one to guess and check. i assume there is something else like "consecutive'

OpenStudy (anonymous):

1 + 1 +115 works

OpenStudy (anonymous):

oh yeah sorry.. :/

OpenStudy (anonymous):

so lets try 25 + 26 + 27

OpenStudy (anonymous):

too small

OpenStudy (anonymous):

yeah thats wrong

OpenStudy (anonymous):

35 + 36 + 37?

OpenStudy (anonymous):

still too small, but not by much

OpenStudy (anonymous):

38+39+40?

OpenStudy (anonymous):

Lol i feel like im cheating by doing that... :P

OpenStudy (anonymous):

bingo. now for the algebra. if you put one number as n, the the other two are n +1 and n+2 then put \[n+n+1+n+2=117\] \[3n+3=117\] etc

OpenStudy (anonymous):

nice to know the answer in advance. then we can make sure the algebra is right \[3n+3=117\] \[3n=114\] \[n=114\div 3\] whatever that is

OpenStudy (anonymous):

alright i have another for you... im good at the algebra part its just the unecessary words confuse me.. :/

OpenStudy (anonymous):

do i get to guess?

OpenStudy (anonymous):

haha sure.. :P

OpenStudy (anonymous):

One number is 800 less then 3 times the other and there sum is 1,200 find the numbers

OpenStudy (anonymous):

grrr

OpenStudy (anonymous):

ok i guess i can guess lets try 500 and 1500-800

OpenStudy (anonymous):

wow got it on the first try!!

OpenStudy (anonymous):

Nice job! :P

OpenStudy (anonymous):

now algebra... one number is n, other is 3n-800

OpenStudy (anonymous):

i would give you another medal if i could.. :P

OpenStudy (anonymous):

now i am sure you can do the algebra for this one \[n + 3n-800=1200\] etc

OpenStudy (anonymous):

actually the point of guessing is to make sure that the algebra makes sense

OpenStudy (anonymous):

okie dokie.. :D

OpenStudy (anonymous):

i guess 500 where does the next number come from? i multiply by 3 and subtract 800

OpenStudy (anonymous):

so if i guess n, the next number is 3n - 800 and that is what i have \[n+3n-800=1200\]

OpenStudy (anonymous):

ohh see i took 800 from 1200 and etc

OpenStudy (anonymous):

so you solve via \[4n-800=1200\] \[4n=1200+800=2000\] \[n=2000\div 4=500\]

OpenStudy (anonymous):

okay now you confuddled me....